OpenStudy (openstudier):

WILL FAN AND MEDAL. HELP, ASAP, PLEASE. Point G(−7, 4) is translated using the rule (x+10, y−6). What is the x-coordinate of G′ ?'

Mehek (mehek14):

Add 10 to x and subtract 6 from y -7 + 10 = 4 - 6 =